OpenStudy (anonymous):

How many mph is a car traveling if the wheels have a diameter of 3ft and are turning 600 times a minute

OpenStudy (amistre64):

you will have to know how many feet in a mile ... its about 5000 if your roman. and how many minutes in an hour

OpenStudy (amistre64):

\[\frac {a~ft}{b~min}*\frac{n~min}{1~hr}*\frac{1~mile}{m~ft}\]

OpenStudy (amistre64):

spose we would need to calculate feet per minute in there someplace, but thats just a calculation of circumference